# LEGO: Learning and Graph-Optimized Modular Tracker for Online Multi-Object Tracking with Point Clouds

### v2

@@ I Introduction
The paper presents several significant contributions, which are outlined as follows:
  • A novel learning and graph-optimized (LEGO) modular tracker is proposed for online MOT. The core components of this system include a self-attention based GNN and Kalman filter, which are easily comprehensible, tunable, and adaptable to different sensor modalities.
  • An innovative offset correction module is designed to predict the offset between these detection results and the ground truth. The primary purpose of this mechanism is to rectify prediction outcomes, thereby improving the accuracy of the model’s predictive capabilities.
  • The graph dual attention network (GDAN) is introduced to process the bipartite graph input to reduce the information redundancy, and directly generate the score matrix. In addition, a new method is suggested for calculating the final association cost matrix, by integrating the geometry and motion cost matrix with the GDAN score matrix derived from the GDAN module.
@@ III-B Motion Prediction Module
$\bm{SA}=\begin{bmatrix}\bm{I}_{3\times 3}&\delta\bm{I}_{3\times 3}&\frac{1}{2}\delta^{2}\bm{I}_{3\times 3}&\\ \bm{O}_{3\times 3}&\bm{I}_{3\times 3}&\delta\bm{I}_{3\times 3}&\bm{O}_{9\times n}\\ \bm{O}_{3\times 3}&\bm{O}_{3\times 3}&\bm{I}_{3\times 3}&\\ &\bm{O}_{n\times 9}&&\bm{I}_{n\times n}\end{bmatrix},$ (3)
@@ III-E GDAN Score Calculation Module
Utilizing the feature map extracted from the feature extraction module, a GDAN score matrix, denoted as $\bm{B}$, is calculated through the GDAN score calculation module. This calculation takes into consideration the relationships between point cloud voxels of various objects. The GDAN score calculation operates in two key stages. Initially, it constructs a bipartite graph comprising both detection nodes and prediction nodes, which is represented as an adjacency matrix of dimensions $J\times U$, where $J$ represents the number of detection nodes and $U$ denotes the number of prediction nodes. Compared with a traditional graph whose dimension is $\left(J+U\right)\times\left(J+U\right)$, which entails redundant computations due to the simultaneous inclusion of detection and prediction nodes. The bipartite graph mitigate this computational redundancy, effectively reducing the dimension of the association score matrix to $J\times U$. Consequently, the association between predicted and detected objects is framed as a maximum-weighted bipartite matching problem, wherein each edge in the graph corresponds to an association score between two distinct nodes. Formally, the bipartite graph is represented as $\bm{G}=\langle\bm{N},\bm{E}\rangle$, with $\bm{N}$ consisting of two node sets, $\bm{N}_{det}$ (the set of nodes representing detected objects) and $\bm{N}_{pre}$ (the set of nodes representing predicted objects). The cardinalities of these sets are $\left|\bm{N}_{det}\right|$ = $J$ and $\left|\mathbf{N}_{pre}\right|=U$, respectively. The set of edges, $\bm{E}$, captures the associations between nodes, and each edge within $\bm{E}$ encodes an association score representing the strength of the relationship between the corresponding nodes. To initialize the bipartite graph, all edge values are assigned an initial score of 1.
@@
The method in [57] directly applies polynomials to update the feature information on the node through message passing. This could result in redundant information for each node since different nodes should contribute different useful information instead of having the same contribution for message passing. To overcome this limitation, the GDAN module is proposed where an MSA encoder is initially employed to further distil the feature representation of each node. Following this, a self-attention-based mechanism is incorporated to manage message passing and updating within the bipartite graph. Our proposed self-attention blocks are specifically tailored to suit the bipartite graph structure, as depicted in Fig. 5.
@@
$Attention\left(\bm{Q},\bm{K},\bm{V},\bm{B}_{init}\right)=f\left(\bm{Q},\bm{K},\bm{V}\right)\times\bm{W}\times\bm{B}_{init}^{T},$ (6)
@@
This formula represents a cross-entropy loss that measures the difference between the predicted association scores ($B_{j,u}$) and the ground truth of the matching pair ($y_{j,u}$). Beware $y_{j,u}$ is a binary indicator which is either 00 or $1$ as it represents the ground truth of the match between the $u$-th predicted object and the $j$-th detected object. When there is a clear match, $y_{j,u}$ is assigned a value of 1, and 0 otherwise. Due to the absence of ground-truth annotations for each pair, we define a match based on object identity continuity between frames: a pair is considered matched if the closest ground-truth box to the tracking box in the previous frame and the closest ground-truth box to the detection box in the current frame have the same index and their IoU with their closest ground-truth boxes are more than 0.7. The value of $B_{j,u}$ is obtained from the GDAN score matrix $\bm{B}$ and ranges between 0 and 1.
@@ III-F Association Module
In 3D MOT, the data association module plays a crucial role in determining the correspondence between detected and predicted objects, to update the state of the objects. Most of the methods used in 3D MOT rely on the minimization of the total association cost matrix. This is done by evaluating the overlap or distance between the 3D bounding boxes of the predicted and detected objects, using the global nearest neighbor principle and typically the Hungarian algorithm [59].
To improve the existing 3D-based method, an approach that includes angle difference and velocity difference measures is proposed. This approach helps in calculating a comprehensive geometry and motion cost matrix, which is expressed by the mathematical formulation
@@
The $\left(u,j\right)$-th entry in the motion cost matrix $\bm{Mo}$ in Eq. (10) is computed as

### v3

@@ I Introduction
The contributions of this paper are outlined as follows:
  • A learning and graph-optimized (LEGO) modular tracker is proposed for online MOT. The tracker is built upon recent model-based tracker, i.e. PC3T [38], with introducing two additional learning modules, an offset correction module and a self-attention based GNN, which are easily comprehensible, tunable, and adaptable to different sensor modalities.
  • The offset correction module is designed to predict the offset between these detection results and the ground truth. The primary purpose of this mechanism is to rectify prediction outcomes, thereby improving the accuracy of the model’s predictive capabilities. The graph dual attention network (GDAN) is introduced to process the bipartite graph input to reduce the information redundancy, and directly generate the association score matrix.
@@ III-E GDAN Score Calculation Module
Utilizing the feature map extracted from the feature extraction module, a GDAN score matrix, denoted as $\bm{B}$, is calculated through the GDAN score calculation module. Similar to [40], the purpose of employing such matrix is to extract the similarities between different detections and predicted tracks. This calculation takes into consideration the relationships between point cloud voxels of various objects. We first construct the graph. Most of the recent works which applied GNN for 3D MOT, e.g. PTP [53], Rethinking3DMOT [37] , 3Dmotformer [54], etc, use the fully connected graph with dimension of $\left(J+U\right)\times\left(J+U\right)$. However, such graph entails redundant computations due to the simultaneous inclusion of detection and prediction nodes. To this end, we follow GNN3DMOT [43] to use the prior knowledge that the matching should only happen across frames that we construct a bipartite graph. The structure of the graph is represented as a matrix of dimensions $J\times U$, where $J$ represents the number of detection nodes and $U$ denotes the number of prediction nodes. The bipartite graph mitigate this computational redundancy, effectively reducing the dimension of the association score matrix to $J\times U$. Consequently, the association between predicted and detected objects is framed as a maximum-weighted bipartite matching problem, wherein each edge in the graph corresponds to an association score between two distinct nodes. Formally, the bipartite graph is represented as $\bm{G}=\langle\bm{N},\bm{E}\rangle$, with $\bm{N}$ consisting of two node sets, $\bm{N}_{det}$ (the set of nodes representing detected objects) and $\bm{N}_{pre}$ (the set of nodes representing predicted objects). The Cardinalities of these sets are $\left|\bm{N}_{det}\right|$ = $J$ and $\left|\mathbf{N}_{pre}\right|=U$, respectively. The set of edges, $\bm{E}$, captures the associations between nodes, and each edge within $\bm{E}$ encodes an association score representing the strength of the relationship between the corresponding nodes. To initialize the bipartite graph, all edge values are assigned an initial score of 1.
@@
The method in [59] directly applies polynomials to update the feature information on the node through message passing. This could result in redundant information for each node since different nodes should contribute different useful information instead of having the same contribution for message passing. To overcome the limitations in [59], GNN3DMOT [43] use graph attention network (GAT) [60] to realize the message passing and updating. However, GAT belongs to the spatial convolution and process the message passing and updating node by node which ignoring the whole structure of the graph. This could miss the important information in the graph we construct. To overcome these limitations, we proposed the GDAN module. We first use an MSA encoder to simple extract the feature representation of each node. Considering the input are point cloud in two sequence frames, we also design an embedding network with LSTM and several fully connect layers to further extract the embedding representation of features, then a modified self-attention-based mechanism is incorporated to manage message passing and updating within the bipartite graph. Different from directly using self-attention like GAT, we add a learnable adjacency matrix into the calculation of the self-attention during the message passing and updating process which can provide the structure information. Our proposed modified self-attention blocks are specifically tailored to suit the bipartite graph structure, as depicted in Fig. 5.
@@
$Attention\left(\bm{Q},\bm{K},\bm{V},\bm{B}\right)=f\left(\bm{Q},\bm{K},\bm{V}\right)\times\bm{W}\times\bm{B}^{T},$ (5)
@@
This formula represents a cross-entropy loss that measures the difference between the predicted association scores ($B_{j,u}$) and the ground truth of the matching pair ($y_{j,u}$). Beware $y_{j,u}$ is a binary indicator which is either 00 or $1$ as it represents the ground truth of the match between the $u$-th predicted object and the $j$-th detected object. When there is a clear match, $y_{j,u}$ is assigned as 1, and 0 otherwise. Due to the absence of ground-truth annotations for each pair, we define a match based on object identity continuity between time frames: a pair is considered as ”matched” if the closest ground-truth box to the tracking box in the previous time frame and the closest ground-truth box to the detection box in the current time frame have the same index and their IoU with their closest ground-truth boxes are more than 0.7. The value of $B_{j,u}$ is obtained from the GDAN score matrix $\bm{B}$ and ranges between 0 and 1.
@@ III-F Association Module
In 3D MOT, the data association module plays a crucial role in determining the correspondence between detected and predicted objects, to update the state of the objects. Most of the methods used in 3D MOT rely on the minimization of the geometry association cost matrix. This is done by evaluating the overlap or distance between the 3D bounding boxes of the predicted and detected objects, using the global nearest neighbor principle and typically the Hungarian algorithm [61]. Recently, [38] proposed to incorporate geometry, motion and appearance features from images to construct such association cost matrix. In our proposed approach, we follow the similar idea to define the geometry and motion cost matrix, but using GDAN score matrix $\bm{B}$ to replace the appearance features. The details are discussed in the follow parts.
Same as [38], the combination of geometry and motion cost matrix is expressed by the mathematical formulation